Circular No. 8612

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V476 SCUTI = NOVA SCUTI 2005 N. N. Samus, Russian Academy of Sciences, informs us that the des-

ignation V476 Sct has been given to this nova (cf. IAUC 8607).

L. Kiss, University of Sydney; M. Bessel, Australian National University (ANU); and A. Retter, Pennsylvania State University, report that a low-resolution spectrogram (0.27 nm/pixel, range 550–1050 nm), taken with the ANU 2.3-m telescope (+ Double-Beam Spectrograph) at Siding Spring Observatory on Oct. 6.36 UT, shows that V476 Sct is a classical nova. Strong emission lines of hydrogen, oxygen, calcium, magnesium, carbon, and iron can be identified. The prominent Hα emission line has a symmetric profile with FWZI exceeding 4000 km/s. The strongest emission lines include Hα, the O I 777.3-, 844.6-, and 926.4-nm lines, members of the hydrogen Paschen series, the infrared Ca triplet, and the C I 940.6nm line, all with FWZI > 2000 km/s. The overall spectral appearance is similar to that of V443 Sct (N Sct 1989), two weeks after its maximum (Williams et al. 1991, Ap.J. 376, 721).

Visual magnitude estimates: Oct. 1.208 UT, 11.8 (M. Linnolt, Woodside, CA); 1.812, 11.5 (K. Hornoch, Lelekovice, Czech Republic); 1.91, 11.9 (J. M. San Juan, Avila, Spain); 2.82, 11.5 (J. Carvajal, Madrid, Spain); 3.82, 11.6 (Carvajal).

COMET C/2005 S1 (SOHO) Further to IAUC 8604, K. Battams reports that additional images of

this comet show that a peak magnitude of 3.4 was reached at ∼ 10 R_{⊙}. The peak tail length in C2 images was ∼ 1^{o}; by the time that the comet disappeared behind the C2 occulter its tail was the same width as its head (so, in essence, it was headless). The final astrometry and improved orbital elements for this Kreutz sungrazer appear on MPEC 2005-T35.

COMET P/2005 S3 (READ) This comet (cf. IAUC 8608) appears to be of short period; the following

improved orbital elements are taken from MPC 2005-T27:

T = 2005 Oct. 9.433 TT  ω = 115.^{o}784
e = 0.43617  Ω = 276.022 }2000.0
q = 2.86795 AU  i = 3.375
a = 5.08659 AU n^{o} = 0.085914 P = 11.47 years
2005 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green